# Improving Clinical Note Generation from Complex Doctor-Patient Conversation

### v1

@@ Introduction
Keyword Extraction
Besides generating SOAP content, we further generate the keyword section for each clinical dialogue to reduce the reviewing complexity. We treat this task as a NER task but with relation as the prefix. For keyword extraction, we use Label-supervised unmasked Llama finetuning (Li et al. 2023). Instead of autoregressive generation, we add a label predictor at the top of a Llama-7b model. The label predictor is a one-layer linear feed forward neural network. The dialogue sequences decoded by Llama will be fed into this predictor to predict the labels. We also keep the idea of removing the causal masks because their empirical studies show that using token representations learned with causal masks significantly underperforms in token classification tasks. In order to make this task more understandable to the model, we merge the entities into only 4 types, which are PRESENT, ABSENT, FAMILY and UNKNOWN. To be specific, we set the word to tag dictionary as ‘O’: 0, ‘B-PRESENT’: 1, ‘B-ABSENT’: 2, ‘B-FAMILY’: 3, ‘B-UNKNOWN’: 4, ‘I-PRESENT’: 5, ‘I-ABSENT’:6, ‘I-FAMILY’:7, ‘I-UNKNOWN’:8 while B means start and I means inside. The model will determine whether a word is a symptom or disease and what is the relation with the patient at the same time.
Evaluation of K-SOAP Note Generation
We present various evaluation metrics on the benchmarking of CliniKnote, to investigate the generated notes’ quality and time complexity reduction. We aim to demonstrate the effectiveness of CliniKnote in producing high-quality clinical notes and its potential to enhance the workflow in medical practice by saving time and improving accuracy.
Evaluation Metrics
For keyword evaluation, we only record and compare the accuracy, precision, recall and F1 score. For SOAP part note evaluation, we employ several metrics to ensure a comprehensive assessment. We start with the most traditional metric, ROUGE (Lin 2004), which evaluates similarity based on common subsequences between the generated and reference texts. In addition to ROUGE, we introduce more advanced metrics: BERTScore (Zhang* et al. 2020) and BLEURT  (Sellam, Das, and Parikh 2020)  (Pu et al. 2021). These metrics leverage contextual embeddings from BERT to evaluate the semantic similarity between generated and reference texts. BERTScore uses token embeddings to compute precision, recall, and F1 scores, providing a more nuanced and context-aware evaluation compared to traditional metrics. Similarly, BLEURT incorporates pre-trained transformer models to better capture the meaning and fluency of the text, offering insights into the semantic quality of the generated notes. We use Roberta-large model for BERTScore metric and the newest checkpoint BLEURT-20 for BLEURT metric.
Finally, we include QuestEval (Scialom et al. 2021) in our evaluation framework. Unlike traditional metrics, QuestEval does not require any ground truth references. Compared with ROUGE, QuestEval substantially improves the correlation with human judgments over four evaluation dimensions: consistency, coherence, fluency, and relevance. By incorporating question-answering models, QuestEval evaluates the generated text’s ability to provide accurate and relevant information, making it particularly suitable for complex domains like clinical note generation.
Model Acc F1 Pre Recall Eval_Loss
biobert 0.933 0.735 0.755 0.715 0.197
scibert 0.921 0.669 0.723 0.622 0.221
pubmedbert 0.919 0.685 0.689 0.681 0.230
bluebert 0.903 0.554 0.679 0.467 0.285
bert 0.915 0.653 0.677 0.630 0.242
Llama2-7b-in-context - 0.692 0.391 0.500 -
Llama2-7b-QA - 0.642 0.731 0.657 0.307
Llama2-7b-
label-supervised 0.971 0.752 0.726 0.780 0.107
Table 4: Keyword performance comparison of different models.
@@
1 ”Chief Complaint”: ”Skin rash on arms and legs.” ”Chief Complaint”: ”Skin issues on forearm.”
2 ”History Of Presenting Illness”: ”62-year-old female presents with a skin issue that started as a small red patch on her forearm about a month ago. The patch has been getting larger and itchier, rated as a 7 out of 10 in severity. Cold compresses provide some relief, while heat and wearing long sleeves aggravate the condition. The itching is affecting her sleep and self-esteem.” ”History Of Presenting Illness”: ”62-year-old female presents with a skin issue that started about a month ago. Initially a small red patch on the forearm, which has been getting larger and itchier. Itchiness rates at about a 7 out of 10. Symptoms worsen at night and when wearing long sleeves. Cold compresses provide slight relief, while heat and fabric contact aggravate the condition. The issue has impacted sleep and caused self-consciousness about appearance. No prior occurrence of similar symptoms. Over-the-counter creams have been tried with little effect. ”
@@
5 ”Social History”: ”Patient is a retired librarian, widowed with two adult children. She enjoys gardening and has never smoked. Drinks wine occasionally with dinner.” ”Social History”: ”Retired librarian, widowed with two adult children. Enjoys gardening, non-smoker, minimal alcohol consumption.”
@@
Experiments
Experimental Settings
We train all the models mentioned in this benchmarking with one A100-80GB GPU. For clinical note generation task, we train all the large language models for two epochs. We use the default setting for question-answer finetuning with QLoRA. We set the learning rate to 0.0002 and the warmup ratio to 0.03. For QLoRA hyperparameters, we set LoRA $r$ to 64 and the loRA $\alpha$ to 16 and all the models are quantized in 4 bits. For keyword task, we just use LoRA without quantizing the model. We set batch size = 8, learning rate = 1e-4, Max length = 256, LoRA $r$ = 256, LoRA $\alpha$=768, LoRA dropout=0.1. We train Llama2-7b model for 5 epochs.
Clinical Note Generation Results
In this section, all models are finetuned and tested in 3 modes, which are full note mode, section-4 mode and section-15 mode that contain different numbers of LoRA adapter (Hu et al. 2021). Full note mode means we only train one adapter to generation the full clinical note. Section-4 mode means we train four different adapters to generate the full note. How we divide the four parts is a little different from how we separate Subjective, Objective, Assessment, and Plan parts. We only have Chief Complaint and History Of Presenting Illness in the first part. The second part contains Past Medical History, Past Surgical History, Family History, Allergies, Social History, Medication List, Immunization History and Review of Systems. The third part contains Vital Signs and Physical Exam while the final part contains Assessment, Diagnosis and Plans. Section-15 mode means we train 15 different adapters to generate all the fifteen sections contained in the note separately. As for commercial models, we try full note zeroshot generation and oneshot(one demo in prompt) in-context learning.
For baseline models, we choose Bart and Biobert as baseline models. Bidirectional autoregressive transformer Bart is the SOTA transformer model for note generation. For matching clinical tasks, we use the continued pre-training on PubMed abstract model of BART-large version. We further compare the performance of BioBart, which has the same model structure as BART but possesses distinct tokenizers and vocab- ulary size.
Keywords [‘1rash’, ‘—’, ‘1itchier’, ‘5at’, ‘5night’, ‘—’, ‘5sweat’]
[‘1rash’, ‘—’, ‘1itchier’, ‘5at’, ‘5night’]
Keywords [‘2cough’, ‘—’, ‘2shortness’, ‘6of’, ‘6breath’, ‘—’, ‘2changes’, ‘6in’, ‘6urination’, ‘—’, ‘2blood’, ‘6in’, ‘6the’, ‘6urine’, ‘—’, ‘2changes’, ‘6in’, ‘6bowel’, ‘6movements’, ‘—’, ‘2nausea’, ‘—’, ‘2vomiting’, ‘—’, ‘2abdominal’, ‘6pain’]
[‘2cough’, ‘—’, ‘2shortness’, ‘6of’, ‘6breath’, ‘—’, ‘2changes’, ‘6in’, ‘6urination’, ‘—’, ‘2blood’, ‘6in’, ‘6the’, ‘6urine’, ‘—’, ‘2changes’, ‘6in’, ‘6bowel’, ‘6movements’, ‘—’, ‘2nausea’, ‘—’, ‘2vomiting’, ‘—’, ‘2abdominal’, ‘6pain’]
Table 3 shows the performance of all the models with the automatic metrics we pick. qCammel-13b-section-15 achieves overall best performance on ROUGE and BERTScores among all open-scouce models while GPT-4o-oneshot is considered the best among all commercial models. To our surprise, section-15 mode does not show significant enhancement on the performance even though it takes 11 and 14 more adapters than other modes. All three modes of Llama3-8b show decent results. The section-4 mode of Llama3-8b shows second best peformance on ROUGE score and full mode has highest score on Bleurt, QuestEval and second highest score on Bertscore among all open-source models. However, OpenBioLLM-Llama3-8b has poor performance on almost all metrics. All three modes of OpenBioLLM-Llama3-8b shows lowest scores on ROUGE, BertScore, Bleurt and QuestEval, even lower than baseline models. The domain knowledge insertion triggers decreasing on information extraction ability. The benchmarking shows that appropriate in-domain instruction is beneficial while it could also jeopardize the model’s original extraction ability. Table 5 shows some sections generated by qCammel-13b-section-15 and compares them with the reference sections. After data augmenting, the occurrence rate of penicillin as an allergy drops by 11.5% The actual occurrence rate of penicillin is 19.5% in test set. However, the qCammel-13b-section-15 model predicts penicillin as an allergy 31.5% of the time without data augmentation. Specifically, the model tends to output penicillin when no allergy is mentioned in the dialogue. After applying data augmentation, the penicillin prediction rate drops to 20%, which aligns much closer to the real distribution.
Keyword Extraction Results
Conclusion
@@
In this study, we addressed the critical task of automatic clinical note generation to enhance the efficiency and accuracy of medical documentation. We introduced the K-SOAP format, which includes keywords at the top of traditional SOAP notes for quicker review and comprehension. Additionally, we presented CliniKnote, a large-scale dataset containing complex dialogues paired with K-SOAP notes. Training large language models on CliniKnote demonstrated substantial improvements in generating accurate clinical notes. Our comprehensive evaluation ensured a thorough assessment of the generated notes, making the assessment of the quality of generated notes more straightforward and robust. This work advances clinical note generation, driving further research and saving doctors time in documentation tasks.
Limitations
Although CliniKnote is created by experts to simulate real-life clinical inquiries, there are still robustness issues as 100% of the data in our dataset are clean. In real-life conversations, ASR systems may not always produce clean transcripts due to various factors such as accents, interjections, and expression variations. These factors pose significant challenges for the robustness of clinical note generation.
Additionally, the range of models used in this benchmarking is relatively narrow. Many open-source models with larger sizes or better configurations may be suitable for generating K-SOAP notes, indicating significant potential for improvement in this pipeline.
Finally, the current evaluation metrics do not perfectly reflect the quality of clinical notes, as they struggle to capture key information and crucial errors in the medical domain. We hope that more advanced evaluation metrics for text summarization and note generation will be developed to address this issue. Moreover, human-expert evaluation still need to be completed. We are planning a more extensive evaluation for both CliniKnote and real-life clinical conversations with finetuned models to assess accuracy and the time savings for doctors in real clinical scenarios.
Ethics Statement
No protected health information were used in the creation of this dataset.
Appendix A Example Appendix
K-SOAP Demo
Dialogue D: Good day! I’m Dr. Keir MacDonald, and I’ll be looking after your health concerns today. I see here you’re Mr. Callahan MacLean, correct? How are you feeling this morning?
P: Morning, Dr. MacDonald. I’ve been better, to be honest. I’ve been having some trouble with frequent urination and a burning sensation when I do. It’s been going on for a few days now.
D: I’m sorry to hear that, Mr. MacLean. Can you tell me more about the onset of these symptoms? Have they been constant or intermittent? And on a scale from 1 to 10, how would you rate any pain associated with urination?
P: It started suddenly about four days ago. The discomfort is pretty constant, especially the urge to urinate. The pain is around a 4, but it’s more the frequency that’s troubling me. It’s affecting my work since I have to run to the bathroom so often.
D: Understood. Have you experienced these symptoms before, or is this the first time?
P: This is the first time, and I haven’t had any tests done for it yet.
D: Let’s delve into your medical history a bit. Aside from this, any other chronic conditions or past illnesses I should be aware of?
P: I’ve been managing type 2 diabetes for the past decade, and I have a history of kidney stones.
D: Any allergies, Mr. MacLean?
P: Just penicillin.
D: Are you up to date with your vaccinations, including your flu shot?
P: I am. Got my flu shot last month.
D: How about family medical history? Any similar issues or other conditions in your family?
P: My father had issues with his prostate, and my mother has a history of urinary tract infections.
D: And what about your personal life? Do you live with anyone, are you currently working, and what do you do for leisure?
P: I’m divorced, live alone, and I’m a retired librarian. I enjoy reading and bird watching. I don’t smoke and I’ll have the occasional beer, maybe once a month.
D: Let’s talk about your current medications. What are you taking, and what dosages?
P: I’m on metformin 500 mg twice daily for my diabetes, and that’s about it.
D: Now, I’ll go through a review of systems. Any fever, chills, or recent weight loss?
P: No fever or chills, and my weight has been stable.
D: Any congestion, sore throat, or ear pain?
P: No, nothing like that.
D: Chest pain or palpitations?
P: No, my chest’s been fine.
D: Cough or difficulty breathing?
P: No respiratory issues.
D: Any changes in bowel movements or abdominal pain?
P: No, my bowels have been normal.
D: Dizziness, headaches, or memory issues?
P: No neurological symptoms.
D: Joint pains or stiffness?
P: My knees ache occasionally, but nothing significant.
D: Any intolerance to heat or cold?
P: No, I’m comfortable with the temperature.
D: I see. Let’s proceed with the physical exam. You don’t appear to be in any respiratory distress. Let’s check your heart rate… It’s 78 beats per minute. Blood pressure is 130/85, and respiratory rate is 14 breaths per minute. All within normal ranges.
D: Your affect is calm and you’re well-oriented. No redness or swelling in the head and neck region. Pupils are equal and reactive to light. Heart sounds reveal normal S1 and S2 without murmurs. Lungs have good air entry bilaterally, no wheezing or other sounds. Abdomen is soft, non-tender, with normoactive bowel sounds. Cognitive screening is normal.
P: That’s good to hear, doctor.
D: Based on your symptoms and history, it seems like you may have a urinary tract infection. We’ll need to run a urinalysis to confirm and possibly a culture to identify the bacteria. I’ll prescribe a course of antibiotics that’s safe for someone with a penicillin allergy. We should also monitor your blood sugar closely, as infections can affect glucose levels.
P: Sounds like a plan, Dr. MacDonald.
D: I’ll give you a requisition for the lab work. Try to get it done as soon as possible. In the meantime, increase your water intake to help flush out your system. I’d like to see you back here in a week to review the lab results and check on your progress. If your symptoms worsen or you develop a fever, please go to the emergency room immediately. Do you have any other concerns or questions for me?
P: No, that covers everything. Thank you for your help.
D: You’re welcome, Mr. MacLean. Take care and I’ll see you soon.
Note: Figure LABEL:fig:full_demo shows the corresponding KSOAP note.
Allgeries Augmenting Prompt
allergens = [ ”Milk”, ”Eggs”, ”Fish”, ”Shellfish”, ”Tree nuts”, ”Peanuts”, ”Wheat”, ”Soybeans”, ”Sesame seeds”, ”Mustard”, ”Sulfites”, ”Lupin”, ”Almonds”, ”Cashews”, ”Walnuts”, ”Pistachios”, ”Hazelnuts”, ”Brazil nuts”, ”Pecans”, ”Macadamia nuts”, ”Pine nuts”, ”Gluten”, ”Oats”, ”Rye”, ”Barley”, ”Corn”, ”Kiwi”, ”Banana”, ”Avocado”, ”Strawberry”, ”Citrus fruits”, ”Apples”, ”Cherries”, ”Peaches”, ”Tomatoes”, ”Carrots”, ”Celery”, ”Garlic”, ”Onions”, ”Chili peppers”, ”Peppercorns”, ”Cinnamon”, ”Cloves”, ”Vanilla”, ”Ginger”, ”Anise seeds”, ”Fenugreek”, ”Sunflower seeds”, ”Poppy seeds”, ”Mollusks”, ”Crustaceans”, ”Artificial colors”, ”Artificial flavors”, ”Monosodium glutamate (MSG)”, ”Aspartame”, ”Nutrasweet”, ”Saccharin”, ”Soy lecithin”, ”Soy protein”, ”Whey protein”, ”Casein”, ”Yeast”, ”Tapioca”, ”Seaweed”, ”Agar agar”, ”Mango”, ”Penicillin”, ”Amoxicillin”, ”Cephalexin”, ”Clindamycin”, ”Vancomycin”, ”Sulfa drugs (Sulfonamides)”, ”Tetracycline”, ”Doxycycline”, ”Aspirin” ]
Prompt = You are a data augmenting assistant and you job is to replace the allergies in the following dialogue between a patient and a doctor. The doctor may ask the patient’s allergies and you should change the patient’s answer with one or more random allergens. If the doctor didn’t ask about the allergies, just output the original dialogue and don’t modify anything. The allergies you need to replace with is ”’+allergens+”’ Please organize the output in json format:{{’output’: new dialogue}} or {’output’: ’not modified’} if doctor doesn’t ask in the dialogue.
Dialogue:”’+dialogue

### v2

@@ 6.3 Clinical Note Generation Results
We choose Bart and Biobert as baseline models.For matching clinical tasks, we use the continued pre-training on PubMed abstract model of BART-large version. We further compare the performance of BioBart, which has the same model structure as BART but possesses distinct tokenizers and vocabulary size.
Table 5 shows the performance of all the models with the automatic metrics we pick. qCammel-13b-section-15 achieves overall best performance on ROUGE and BERTScores among all open-scouce models while GPT-4o-oneshot is considered the best among all commercial models. Section-15 mode does not show significant enhancement on the performance even though it takes 11 and 14 more adapters than other modes. All modes of Llama3-8b show decent results. The section-4 mode of Llama3-8b shows second best peformance on ROUGE score and full mode has highest score on Bleurt, QuestEval and second highest score on Bertscore among all open-source models. However, OpenBioLLM-Llama3-8b has poor performance on almost all metrics. All three modes of OpenBioLLM-Llama3-8b shows lowest scores on ROUGE, BertScore, Bleurt and QuestEval, even lower than baseline models. The domain knowledge insertion triggers decreasing on information extraction ability. The benchmarking shows that appropriate in-domain instruction is beneficial while it could also jeopardize the model’s original extraction ability. Table 6 shows some sections generated by qCammel-13b-section-15 and compares them with the reference sections. After data augmenting, the occurrence rate of penicillin as an allergy drops by 11.5% The actual occurrence rate of penicillin is 19.5% in test set. However, the qCammel-13b-section-15 model predicts penicillin as an allergy 31.5% of the time without data augmentation. Specifically, the model tends to output penicillin when no allergy is mentioned in the dialogue. After applying data augmentation, the penicillin prediction rate drops to 20%, which aligns much closer to the real distribution.
@@
1 "Chief Complaint": "Skin rash on arms and legs." "Chief Complaint": "Skin issues on forearm."
2 "History Of Presenting Illness": "62-year-old female presents with a skin issue that started as a small red patch on her forearm about a month ago. The patch has been getting larger and itchier, rated as a 7 out of 10 in severity. Cold compresses provide some relief, while heat and wearing long sleeves aggravate the condition. The itching is affecting her sleep and self-esteem." "History Of Presenting Illness": "62-year-old female presents with a skin issue that started about a month ago. Initially a small red patch on the forearm, which has been getting larger and itchier. Itchiness rates at about a 7 out of 10. Symptoms worsen at night and when wearing long sleeves. Cold compresses provide slight relief, while heat and fabric contact aggravate the condition. The issue has impacted sleep and caused self-consciousness about appearance. No prior occurrence of similar symptoms. Over-the-counter creams have been tried with little effect. "
@@
5 "Social History": "Patient is a retired librarian, widowed with two adult children. She enjoys gardening and has never smoked. Drinks wine occasionally with dinner." "Social History": "Retired librarian, widowed with two adult children. Enjoys gardening, non-smoker, minimal alcohol consumption."
@@
The inclusion of keyword guidance significantly enhances note generation for complex sections, achieving a 21.5% increase in F1 score for the Assessment section and a 22.5% increase for the History of Presenting Illness section. The addition of keywords helps guide the model to focus more effectively on the correct information. Besides the accuracy improvement of generation, we also notice obvious prevention of hallucinations, as verified through a manual review of the output differences. However, a noticeable performance drop is observed in simpler sections such as Chief Complaint and Allergies. This decline likely occurs because these sections do not rely on keyword information, which may introduce noise rather than value.
6.4 Conclusion
@@
This study introduces automatic clinical note generation to improve medical documentation efficiency and accuracy with the guidance of keywords. We introduce the K-SOAP format, adding keywords to traditional SOAP notes for quicker review, and CliniKnote, a large-scale dataset of complex dialogues paired with K-SOAP notes. Training large language models on CliniKnote significantly improved clinical note accuracy. Our evaluation highlights its potential to streamline documentation and save doctors valuable time.